A student has conducted an experiment to measure the conductivity of electrolyte

Y2SO4 and was found to be 0.147 Ω-1 m-1. This solution was prepared by reacting 23.50

mL of 0.012 mol dm-3 of H2SO4 and YOH that gives the total volume of reaction

mixture is 50 mL.

(a) Write the chemical equation for the reaction.

(b) Find the concentration of YOH

(c) Find the limiting molar conductivity of Y ion


1
Expert's answer
2021-01-05T05:57:09-0500

The chemical equation for the reaction is:

2YOH + H2SO4 "\leftrightarrow" Y2SO4 + 2H2O.

The total volume of reaction mixture is 50 mL, therefore the volume of the YOH reacted is:

"V(YOH) = 50 - 23.50 = 26.50" mL.

As it can be seen from the reaction equation, the volume number of the moles of YOH that react is twice the number of the moles of H2SO4:

"n(YOH) = 2n(H_2SO_4)" .


Therefore, the concentration of YOH is:

"c(YOH) = \frac{n(YOH)}{V(YOH)} = \frac{2n(H_2SO_4)}{V(YOH)}"

"c(YOH) = \frac{2c(H_2SO_4)V(H_2SO_4)}{V(YOH)} = \frac{2·0.012·23.50·10^{-3}}{26.50·10^{-3}}"

"c(YOH) = 0.0213" mol dm-3.


Then, the concentration of Y2SO4 is:

"c(Y_2SO_4) = \frac{c(H_2SO_4)·V(H_2SO_4)}{V_r} = \frac{0.012·23.50·10^{-3}}{50·10^{-3}}"

"c(Y_2SO_4) = 0.00564" mol dm-3.



The molar conductivity of the reaction solution is:

"\lambda_m = \frac{\kappa}{c} = \frac{0.147}{0.00564·10^3} = 0.026" Ω-1 m2 mol-1, or 261 Ω-1 cm2 mol-1.

As the concentrations are small, we can consider this molar conductivity as a limiting molar conductivity.

"\lambda_m\approx\lambda_m^0" .

The limiting conductance of SO42- is 160 Ω-1 cm2 mol-1. Finally, the limiting molar conductivity of Y ion is:

"\lambda_m^0(Y^{+}) =\frac{\lambda_m^0 - \lambda_m^0(SO_4^{2-})}{2} = \frac{261- 160}{2} = 51" Ω-1 cm2 mol-1.


The value resembles to that of sodium ion, Na+.